Father is 53 and son is 17.How many years ago father was 10 times older than his son?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

53-x = 10(17-x)
53-x = 170-10x
9x = 117
x = 13
13 years ago the father was 10 times older than his son
-----------------------------------------------------
check:
53-13 = 40 father
17-13 = 4 son
and 4 times 10 = 40
Correct!

Answer by ikleyn(52781)

.
Father is 53 and son is 17.How many years ago father was 10 times older than his son?
~~~~~~~~~~~~~~~~~~~~~~~~


For your info:


In English,  "father was 10 times older than his son"  means  "father was 11 times as old as his son".